# arXiv:2603.01660v1[eess.SP] 02 Mar 2026

### v1

@@ V Results
We evaluate the performance of the proposed radar–IRS system using a simulated 24 GHz sensing configuration. The radar transmitter (Tx) and receiver (Rx) are co-located at the origin, $(0,0,0)$ m, and operate at a carrier frequency of $f_{0}=24$ GHz, corresponding to a wavelength of $\lambda=0.0125$ m. The receive aperture is a uniform linear array (ULA) composed of $M_{R}=50$ elements placed along the $x$-axis, with half-wavelength spacing $d_{\mathrm{radar}}=\lambda/2=6.25$ mm. Hence, the $m$-th receive sensor is located at $\bigl((m-1)d_{\mathrm{radar}},\,0,\,0\bigr)$.
Figure 3: Proposed Configuration in 3-Dimensional Space
A single IRS panel is positioned at (50, 100, 0)(-50,\,100,\,0) m as shown in Fig. 3 and realized as a $20\times 18$ rectangular grid of reflecting elements, resulting in $L_{\mathrm{total}}=360$ controllable units. The IRS elements are arranged in the yyzz plane with half-wavelength spacing along both axes, i.e., $d_{\mathrm{irs},y}=d_{\mathrm{irs},z}=\lambda/2$. In addition to geometric placement, the complex reflection weights of the IRS are designed to jointly align the incident radar signal from the target direction assumed to be at $\bigl[5,\,35,\,18\bigr]$ with the desired outgoing look-up direction toward the radar receiver. All propagation is modelled under free-space conditions with wave speed $c=3\times 10^{8}$ m/s, and half-wavelength sampling is employed across all arrays to avoid grating lobes and preserve directional integrity at 24  GHz.
Figure 4: RMSE vs SNR behavior of IRS-static Radar
Fig. 4 shows that azimuth and elevation angle RMSEs (on a logarithmic scale) decrease steadily as SNR increases from 15-15 dB to $15$ dB, confirming improved estimation accuracy with stronger signals. The azimuth and elevation curves remain separated across all SNRs, indicating systematically better accuracy in one angular dimension due to the radar–IRS geometry and its effective aperture and directional gain. At moderate-to-high SNRs, the RMSE drops sharply, revealing that reliable sub-degree angle estimation is achievable once the SNR surpasses a mid-range threshold.
Figure 5: RMSE vs number of Snapshots
The Figure 5 analyzes how azimuth and elevation RMSE vary with the number of snapshots (on a logarithmic scale). Both errors decrease as snapshots increase from hundreds to thousands, with the largest RMSE in the low-snapshot regime and diminishing improvements thereafter as the covariance estimate stabilizes. Throughout, elevation and azimuth RMSE curves remain separated, reflecting the anisotropic array–IRS geometry that inherently favors one angular dimension. Overall, the results show that structural aspects of the sensing configuration drive performance disparities, and that increasing snapshots significantly enhances angle estimation reliability, especially at moderate SNR.
@@
Figure 6: RMSE vs Deviation in Azimuth of Target from the IRS Target lookup direction
Figure 7: RMSE vs Deviation in Elevation of Target from the IRS Target lookup direction
In our experiment, we have now set the IRS look-up direction and the nominal target position to $\bigl[5,\,35,\,18\bigr]$. Controlled signed perturbations were applied to the target’s azimuth and elevation angles seperately within $\pm 10^{\circ}$, and the resulting changes in the Cramér–Rao lower bound (CRLB) were analyzed as the target deviated from the look-up direction, as shown in Figures 6 and 7. The RMSE increased with mismatch magnitude, indicating reduced Fisher information due to misalignment between the reflected field and the beam. The symmetry around zero mismatch suggests that equal under- and over-steering have similar effects on accuracy. Increasing the number of snapshots from $10$ to $1000$ improved covariance estimation, lowered the CRLB, and reduced the RMSE across the entire error range.
Figure 8: RMSE vs Number of IRS elements
Figure 8 shows how azimuth and elevation RMSE depend on IRS elements, using a logarithmic scale. As $N_{IRS}$ increases from 0 to $100$, both RMSE values decline steadily, highlighting the impact of aperture gain in IRS-assisted sensing. Larger IRS panels enhance array gain and phase shaping, boosting reflected signals and angular resolution. RMSE reduction is more significant between $N=0$ and $40$, then flattens, indicating diminishing returns as aperture grows. Elevation RMSE always exceeds azimuth RMSE, reflecting anisotropy from the radar–IRS setup and array geometry. Overall, increasing IRS elements benefits the low- to mid-range aperture, but optimising design balances hardware costs and performance at larger sizes.
@@ VI Conclusion
This study examined NLoS sensing and stealth-target detection using an IRS-based configuration termed ‘IRS-static radar’. Unlike traditional IRS-assisted radar that requires a significant four-hop path, this setup uses signals scattered by the target, received at the IRS and sent back to the radar, mimicking bi-static sensing but with a mono-static radar front end. This method offers a practical way to detect low-observable targets. A systematic signal model for the IRS-static radar was developed and reformulated for target parameter estimation. Using this model, we derived closed-form Fisher Information Matrix (FIM) and Cramér-Rao Lower Bounds (CRLB) for target angle estimation, making the analysis more straightforward.
Numerical results demonstrated that estimation accuracy improves with higher SNR and more snapshots. The system’s sensitivity to IRS pointing and beamforming was observed, with performance decreasing as the target moves away from the steering direction. Increasing IRS elements improved bounds due to better passive beamforming and resolution. Elevation estimates were more accurate than azimuth estimates due to the array geometry, guiding the optimization of IRS size and steering for reliable parameter estimation in challenging environments. Overall, the IRS-static radar enhances sensing capabilities and addresses limitations of current IRS-assisted systems. Future work includes optimizing IRS placement and phase profiles, and validating them experimentally with measured channels and hardware prototypes.
In addition to the work carried out, we will develop an estimation algorithm and evaluate the performance of various estimation algorithms for the configuration and signal model discussed.

### v2

@@ V Results
We evaluate the performance of the proposed IRS-static radar system using a simulated 24  GHz sensing configuration. The radar transmitter (Tx) and receiver (Rx) arrays are co-located at the origin of the coordinate system. The transmitter and receiver array apertures are uniform linear arrays (ULA) composing $M_{T}=M_{R}=50$ elements placed along the $x$-axis, with half-wavelength spacing $d_{\mathrm{radar}}=\lambda/2=6.25$ mm. Hence, the $m$-th receive sensor is located at $\bigl((m-1)d_{\mathrm{radar}},\,0,\,0\bigr)$.
A single IRS panel is positioned at $\textbf{r}_{I}=\bigl[-50,\,100,\,0\bigr]^{T}$ m and realized as a $20\times 18$ rectangular grid of reflecting elements, resulting in $L=360$ controllable units. The IRS elements are arranged in the yyzz plane with half-wavelength spacing along both axes. The FOV of the IRS-static radar is the intersection of the transmit beam and the IRS look direction, as shown in Fig. 1(c). Accordingly, the IRS weights are computed using (7) so that the FOV is steered to the point $(\Psi^{(\mathrm{az})}_{0},\Psi^{(\mathrm{el})}_{0})$, and the reflected signal is directed toward the radar receive array at $(\Psi^{(\mathrm{az})}_{\mathrm{R}},\Psi^{(\mathrm{el})}_{\mathrm{R}})$. Being agnostic to any parameter estimation technique, we study the CRB against various system and environment parameters.
((a))
((b))
@@
Figure 2: RMSE performance versus angular mismatch between the target and IRS lookup direction in (a) Azimuth and (b) Elevation.
[…]
It is interesting to note that the RMSE curve for azimuth angle deviation exhibits a narrow beam around $\Psi^{\mathrm{(az)}}_{0}$ as seen in Fig. 2(a), compared to that observed for elevation angle around $\Psi^{\mathrm{(el)}}_{0}$ in Fig. 2(b). This is because of the large effective aperture due to the IRS elements and the radar receive array in the x-y plane. For the elevation angle estimation, the wide RMSE curve around $\Psi^{\mathrm{(el)}}_{0}$ in Fig. 2(b) can be attributed to the reduced effective aperture of IRS elements present along the z-plane. Furthermore, the inverted sidelobes do not exhibit significantly high RMSE compared to the mainlobe, indicating target response from outside the FOV will receive a reasonable gain unless the target lies exactly on the beams’ nulls. From this observation, it can be inferred that the RMSE performance, the FOV, and the sidelobe performance of the IRS-static radar can be altered through appropriate design of the IRS weights. Conversely, the required FOV along the azimuth and elevation angles have to be critical design parameters for IRS weights in this context.
((a))
((b))
Figure 3: RMSE performance versus (a) SNR and (b) Number of snapshots.
[…]
A monotonous CRB decrease is observed in Fig. 3(a) with increasing SNR. For $\Psi_{0}=\Psi_{k}$, the RMSE error for azimuth angle is distinctly lesser than the elevation angle over the entire SNR range due to the larger effective aperture along the x-y dimension over the z-dimension as discussed earlier. Similarly, Fig. 3(b) examines the dependence of azimuth and elevation RMSE on the number of snapshots. Both error metrics decrease as the number of snapshots increases from the order of hundreds to the order of thousands, with progressively diminishing gains thereafter, once the covariance estimate becomes more stable.
When we specifically observe the plots corresponding to deviation of $0.5^{\circ}$, i.e., $\Psi^{(\mathrm{az})}_{0}-\Psi^{(\mathrm{az})}_{k}=0.5^{\circ}$ and $\Psi^{(\mathrm{el})}_{0}-\Psi^{(\mathrm{el})}_{k}=0.5^{\circ}$ (azimuth–pink and elevation–black), one can note that the RMSE is higher than the condition when $\Psi^{(\mathrm{az})}_{0}=\Psi^{(\mathrm{az})}_{k}$ and $\Psi^{(\mathrm{el})}_{0}=\Psi^{(\mathrm{el})}_{k}$ because the target is off the IRS look direction. The azimuth RMSE is lower than elevation RMSE, similar to the previous case, as both these deviations lie in the main lobe corresponding to Fig. 2(a) and Fig. 2(b). For the case where $\Psi^{(\mathrm{az})}_{0}-\Psi^{(\mathrm{az})}_{k}=2.5^{\circ}$ and $\Psi^{(\mathrm{el})}_{0}-\Psi^{(\mathrm{el})}_{k}=2.5^{\circ}$, the corresponding plots (azimuth–green and elevation–gold) in Fig. 3(a) and Fig. 3(b) reveal that the RMSE curves are higher than those of the $0^{\circ}$ and $0.5^{\circ}$ cases. This is due to the deviation from the look-up direction being larger compared to the $0.5^{\circ}$ case, resulting in larger loss in directional gain at the IRS. Furthermore, it is interesting to note that the elevation RMSE is less than the azimuth RMSE. This is because the target azimuth angle lies in the first sidelobe of the pattern observed in Fig. 2(a) while the target elevation angle still resides in the mainlobe pattern shown in Fig. 2(b).
Figure 4: RMSE performance as a function of the number of IRS elements.
Figure 4 illustrates the dependence of azimuth and elevation RMSE on the number of IRS elements, $L=L_{h}L_{v}$. In this study, as the number of IRS elements increases, we maintain a square IRS panel with $L_{h}=L_{v}\in[2,40]$. As $L$ increases from $4$ to $1600$, the azimuth angle and elevation angle RMSE decrease, proving the role of aperture gain in IRS-assisted sensing. Larger IRS apertures enhance the effective array gain and enable more accurate phase synthesis, thereby strengthening the reflected signals and improving angular resolution. The RMSE reduction is most pronounced in the range $4\leq L\leq 400$, after which the curves tend to saturate, indicating diminishing returns with further aperture enlargement. The RMSE of azimuth is lower than that of elevation, following the similar trend observed in Figs. 3(a) and  3(b). Overall, increasing the number of IRS elements is particularly beneficial in the low- to mid-aperture regime, whereas for larger apertures the design must balance incremental performance gains against the FOV and sidelobe attenuation.
Furthermore, for the design of IRS weights as illustrated in (7), it is to be noted that the beamwidth decreases with increasing aperture length, thereby affecting the FOV. We remark that the design of IRS weights using well-constructed optimization problem takes the center stage in the performance of the IRS-static radar configuration for the desired FOV.
@@ VI Conclusion
In this work, a three-hop IRS-assisted radar configuration, termed as IRS-static radar, is considered for study. It is shown that this configuration mimics a bi-static radar with time synchronization alleviated with IRS panel redirecting the scattered signal back to the mono-static radar. A signal model is formulated and compactly represented for this configuration to derive the CRB for target parameters. The dependence of CRB on the system parameters such as the SNR, number of snapshots, number of IRS elements and their weights is brought forward through extensive simulations. The CRB limit is studied as the target moves away from the IRS look direction, both in azimuth and elevation directions, seeking appropriate design for the FOV. It is therefore inferred that the beamformer design for IRS takes center stage in the performance of the IRS-static radar. This study enables a designer to customize the system parameters as per the system specification and serves as a benchmark for parameter estimation techniques developed for this configuration.